### v1

@@ 4.1.3 Planar Boundary Surface Area Test
We repeat the test from 4.1.2 but with two subdomains created using a planar boundary as described in 3.5.2; the boundary curves are shown in Figure 3 (Right), discretized with 2000 points. The $z=0$ plane is used to separate the subdomains. The Hilbert space given by (35) is used again, with $d_{n}$ replacing $d_{x,n_{x}}$, $\left\|\boldsymbol{\omega}_{n}\right\|_{2}$ replacing $\left|\omega_{x,n_{x}}\right|$, and $q=5,T=10$, in a box with dimensions $10\times 6\times 2$ with $\left(2\cdot 13+1\right)^{3}$ Fourier modes. For this test, we use (32) to solve the Laplace-Beltrami problem. The surface area is estimated again in Table 3, with the relative error computed using the $\tilde{N}=64000$ result from Table 2 as an approximate “exact” value. We see that we are able to obtain an accurate estimate with far fewer points than the previous Voronoi test, though the Voronoi test used fewer points per subdomain, which is the main factor in computation time. This method achieves a more accurate estimate using $2560$ points than a triangulation does with 64000, and the 2560 point result is competitive with the 4.8 million vertex triangulation. We also estimate the convergence rate in Table 3 and see that the method appears to converge at a high-order rate as expected.
@@ 4.2 High-Order Approximation of Singular Integrals
Therefore, we need to alter our framework slightly. Rather than a Hilbert space of functions, we consider a Hilbert space $\mathcal{H}$ equipped with a (not necessarily injective) linear evaluation operator $\mathcal{E}$ and consider a problem analogous to (5):
@@
$\mathcal{E}$ in this case must map $a\in\mathcal{H}$ to a function that is $C^{p}$, where $\mathcal{F},\mathcal{G}$ have a maximum order of $p$, except possibly at a finite number of points in $S$ (but not in $S_{\tilde{N}}$ or $\partial S_{\tilde{N}_{\partial}}$). With a suitable choice of $\mathcal{H}$ and $\mathcal{E}$ it is then possible to ensure that the constraint set for (36) is non-empty and closed. If a solution in $a\in\mathcal{H}$ exists such that $\mathcal{F}\left(\mathcal{E}a\right)=f$ on $S$ and $\mathcal{G}\left(\mathcal{E}a\right)=g$ on $\partial S$, convergence is guaranteed due to Proposition 2.1. We illustrate this approach with a couple of examples.
@@ 4.2.1 Singular Integrand in 2D
Integrals of the form $\int_{S}f\left(\boldsymbol{y}\right)K\left(\boldsymbol{x},\boldsymbol{y}\right)\,\text{d}\boldsymbol{y}$, where $K$ has a singularity when $\boldsymbol{x}=\boldsymbol{y}$, are common when solving PDEs by boundary integral approaches and appear in various physical applications, particularly in electromagnetism. Let $S$ be the unit disk and then consider the example:
@@
One possible function $u$ is $\frac{1}{8\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}+\frac{1}{8\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}$; we could simply use this to convert the integral into a boundary integral, but of course the goal is to be able to integrate functions when we do not have an expression for $u$. For example, if $S$ were instead a curved surface and $\Delta u$ were swapped for $\Delta_{S}u$, we would not have such an expression for a possible $u$ available in general.
@@
When $\boldsymbol{x}_{0}\in S$, $\frac{1}{2\pi}\ln\left\|\boldsymbol{x}_{j}-\boldsymbol{x}_{0}\right\|_{2}$ is singular at $\boldsymbol{x}_{0}$, and therefore requires our operator $\mathcal{E}$ to take elements of $\mathcal{H}$ to a function with a singularity at $\boldsymbol{x}_{0}$. To be more concrete, let $\mathcal{H}=\ell^{2}\times\ell^{2}$, then, where $\Omega$ is again a box that contains $S$, let $\mathcal{E}:\mathcal{H}\to C\left(\Omega\right)\cap C^{2}\left(\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}\right)$ be such that
@@
where $u,v\in C^{q}\left(\Omega\right)$ vary with $a$ and $s\in C^{2}\left(\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}\right)$ is fixed and $\Delta s$ has terms that capture the expected singularity. For example, $\Delta\left(u+sv\right)=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ is solved by $u\left(\boldsymbol{x}\right)=\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2},v\left(\boldsymbol{x}\right)=-\frac{1}{16\pi},s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\ln\left(\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\right)$. Assuming we did not know this exact solution, but at least knew $\Delta\left(u+sv\right)$ needed a logarithm singularity with the scaling and any non-singular terms undetermined, we could use $s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2}$ and
@@
where $\left\{\boldsymbol{\omega}_{n}\right\}$ and $\left\{\tilde{\boldsymbol{\omega}}_{n}\right\}$ correspond to Fourier frequencies on the box $\Omega$. Note that $d_{n}$ and $\tilde{d}_{n}$ control the smoothness of $u$ and $v$, respectively.
It is important to note that $\Delta s,\partial_{x}s,$ and $\partial_{y}s$ should not be zero on $\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}$ with a delta distribution singularity at $\boldsymbol{x}$; this would allow arbitrary multiples of $s,xs$ or $ys$ to be added to our solution and change the resulting integral (due to the delta distribution).
For our chosen $s$, a logarithm term appears in $\Delta s$, which will allow $\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ to be integrated. After computing $u+sv$ so that $\Delta\left(u+sv\right)=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ on $\tilde{N}$ scattered points, the boundary integral $\oint_{\partial S}\nabla\left(u+sv\right)\cdot\hat{\boldsymbol{n}}_{\partial S}$ is computed on 1000 evenly spaced points on the unit circle with the standard trapezoid rule (which converges super-algebraically for $C^{\infty}$ functions for the unit circle). We use $d_{n}=\left(\exp\left(q\sqrt{2\pi/T}\right)+\exp\left(q\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2},q=4,T=10$, $\Omega=\left[-2,2\right]^{2}$, and $\left(2\cdot 30+1\right)^{2}$ Fourier basis functions. Interior points are generated using algorithm A.2 of [30] with 20 test points per point and no additional weighting of points near the boundary. The relative error for computing $\frac{1}{2\pi}\int_{S}\ln\left\|\boldsymbol{x}\right\|_{2}\,\text{d}\boldsymbol{x}=\frac{1}{4}$ is shown in Figure 4. High-order convergence is observed, where we note that the point spacing varies as $h_{\text{max}}\propto\tilde{N}^{-\frac{1}{2}}$. We note again that a unit circle is only used as a test problem to compare to a known solution; the method’s main use is for irregularly-shaped domains and surfaces.
@@ 4.2.2 Singular Integrand on a Surface
We now consider the surface described by $S:\left\{\left(x,y,z\right):\varphi\left(x,y,z\right)=0,z\geq 0\right\}$ where $\varphi$ is the level set for a paraboloid: $\varphi\left(x,y,z\right):=z+x^{2}+y^{2}-1.$ We compute
@@
where $\boldsymbol{x}_{0}=\left(0,0,1\right)\in S$. This integral can be interpreted as the electric potential at $\boldsymbol{x}_{0}$ produced by a constant surface charge on $S$. We integrate by attempting to find a function of the form $u+sv$ such that $\Delta_{S}\left(u+sv\right)=-\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$, where $\Delta_{S}$ is now the Laplace-Beltrami operator, and then converting the surface integral to a line integral on the unit circle in the $xy$-plane. The line integral is again evaluated using the trapezoid rule and 1000 evenly spaced points. The Laplace-Beltrami problem is discretized using a method similar to the one used in the previous test (see Eq. (37)), but with $\Delta_{S}$ computed using (31) replacing $\Delta$. We seek a solution of the form $u+sv$ where
$s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2},$
and $u,v\in C^{\infty}\left(\Omega\right)$ are unknown. This choice of $s$ gives $\Delta s=2\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$; however, $\Delta s$ is only one term in $\Delta_{S}s$. The necessary functions $u$ and $v$ to solve $\Delta_{S}\left(u+sv\right)=-\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$ will vary depending on the local geometry of the surface and will be computed numerically. The idea is therefore to use a known Poisson solution in 3D space to handle the desired singular behaviour in the surface Poisson problem.
The Hilbert space and $\mathcal{E}$ used are identical to Subsection 4.2.1, but with $\Omega=\left[-2,2\right]^{3}$ (instead of $\left[-2,2\right]^{2}$) and $\left(2\cdot 11+1\right)^{3}$ Fourier basis functions. Surface points are generated using version B of algorithm A.1 of [30] with 25 test points per point; notably, we do not place extra points near the singularity. This is compared to the naive approach, where a solution $u\in C^{\infty}\left(\Omega\right)$ is sought, in Table 4 (effectively searching for solutions of the form $\mathcal{E}\left(a,0\right)$ only, where $\mathcal{E}$ is similar to (38), but for 3 dimensions). We compare to a value computed to machine precision from a 1D integral resulting from a parametrization, and see that the augmented method converges to the first 9 digits by $\tilde{N}=2560$, while the naive approach fails to converge beyond a digit or two. The Poisson solution using the augmented approach for $\tilde{N}=2560$ is shown in Figure 5.
@@
Figure 5: A computed solution to $\Delta_{S}u=-\frac{1}{4\pi}\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{-1}$ on the paraboloid using the augmented method with $\tilde{N}=2560$ points ($\boldsymbol{x}_{0}=\left(0,0,1\right)$).
@@ 5 Conclusions
We presented, analyzed, and tested two high-order, meshfree techniques for surface integration. The first method (Subsection 3.1) uses a result relating solvability and boundedness (Proposition 2.1) to determine which scaling of a function makes a Poisson problem solvable, then deduces the ratio of the integral of two functions on a domain of interest. This approach is useful for finding the average values of functions on surfaces, as shown in the tests of 4.1.1. We also demonstrated numerically that reasonable estimates of the average value of a function can be obtained using our approach even on irregular, non-uniform point distributions. The second method (Subsection 3.2) uses the divergence theorem to reduce surface integrals to line integrals and takes advantage of the ability of symmetric meshfree methods to find one of many possible solutions to PDEs; this avoids imposing boundary conditions on boundaries that are merely piecewise-smooth. We demonstrated method 2’s accuracy in 4.1.2 and 4.1.3, where we estimated the surface area of a genus-two surface considerably more accurately than a standard triangulation approach. Both methods also preserve the super-algebraic approximation properties of meshfree interpolation methods and do not require prior knowledge of the probability distribution from which a point cloud is sampled; this differs from Monte Carlo methods which are $\mathcal{O}\left(\tilde{N}^{-\frac{1}{2}}\right)$ and typically assume a certain sampling distribution.

### v2

@@ 4.1.3 Planar Boundary Surface Area Test
We repeat the test from 4.1.2 but with two subdomains created using a planar boundary as described in 3.5.2; the boundary curves are shown in Figure 3 (Right), discretized with 2000 points. The $z=0$ plane is used to separate the subdomains. The Hilbert space given by (35) is used again, with $d_{n}$ replacing $d_{x,n_{x}}$, $\|\boldsymbol{\omega}_{n}\|_{2}$ replacing $|\omega_{x,n_{x}}|$, and $q=5,T=10$, in a box with dimensions $10\times 6\times 2$ with $(2\cdot 13+1)^{3}$ Fourier modes. For this test, we use (32) to solve the Laplace-Beltrami problem. The surface area is estimated again in Table 4, with the relative error computed using the $\tilde{N}=64000$ result from Table 3 as an approximate “exact” value. We see that we are able to obtain an accurate estimate with far fewer points than the previous Voronoi test, though the Voronoi test used fewer points per subdomain, which is the main factor in computation time. This method achieves a more accurate estimate using $2560$ points than a triangulation does with 64000, and the 2560 point result is competitive with the 4.8 million vertex triangulation. We also estimate the convergence rate in Table 4 and see that the method appears to converge at a high-order rate as expected.
@@ 4.2 High-Order Approximation of Singular Integrals
Therefore, we need to alter our framework slightly. Rather than a Hilbert space of functions, we consider a Hilbert space $\mathcal{H}$ equipped with a (not necessarily injective) linear operator $\mathcal{E}$ that maps elements of the Hilbert space to a function on $\Omega$. We then consider a problem analogous to (5):
@@
$\mathcal{E}$ in this case must map $a\in\mathcal{H}$ to a function that is $C^{p}$, where $\mathcal{F},\mathcal{G}$ have a maximum order of $p$, except possibly at a finite number of points in $S$ (but not in $S_{\tilde{N}}$ or $\partial S_{\tilde{N}_{\partial}}$). With a suitable choice of $\mathcal{H}$ and $\mathcal{E}$ it is then possible to ensure that the constraint set for (36) is non-empty and closed. If a solution in $a\in\mathcal{H}$ exists such that $\mathcal{F}\,(\mathcal{E}a)=f$ on $S$ and $\mathcal{G}\,(\mathcal{E}a)=g$ on $\partial S$, convergence is guaranteed due to Proposition 2.1. We illustrate this approach with a couple of examples.
@@ 4.2.1 Singular Integrand in 2D
Integrals of the form $\int_{S}f\,(\boldsymbol{y})\,K\,(\boldsymbol{x},\boldsymbol{y})\,\text{d}\boldsymbol{y}$, where $K$ has a singularity when $\boldsymbol{x}=\boldsymbol{y}$, are common when solving PDEs by boundary integral approaches and appear in various physical applications, particularly in electromagnetism. Let $S$ be the unit disk and then consider the example:
@@
One possible function $u$ is $\frac{1}{8\pi}\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{2}\ln\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}+\frac{1}{8\pi}\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{2}$; we could simply use this to convert the integral into a boundary integral, but of course the goal is to be able to integrate functions when we do not have an expression for $u$. For example, if $S$ were instead a curved surface and $\Delta u$ were swapped for $\Delta_{S}u$, we would not have such an expression for a possible $u$ available in general.
@@
When $\boldsymbol{x}_{0}\in S$, $(2\pi)^{-1}\ln\left\|\boldsymbol{x}_{j}-\boldsymbol{x}_{0}\right\|_{2}$ is singular at $\boldsymbol{x}_{0}$, and therefore requires our operator $\mathcal{E}$ to take elements of $\mathcal{H}$ to a function with a singularity at $\boldsymbol{x}_{0}$. To be more concrete, let $\mathcal{H}=\ell^{2}\times\ell^{2}$, then, where $\Omega$ is again a box that contains $S$, let $\mathcal{E}:\mathcal{H}\to C\,(\Omega)\cap C^{2}(\Omega\setminus\{\boldsymbol{x}_{0}\})$ be such that
@@
where $u,v\in C^{q}(\Omega)$ vary with $a$ and $s\in C^{2}(\Omega\setminus\{\boldsymbol{x}_{0}\})$ is fixed and $\Delta s$ has terms that capture the expected singularity. For example, $\Delta\,(u+sv)=-\frac{1}{2\pi}\ln\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}$ is solved by $u\,(\boldsymbol{x})=\frac{1}{8\pi}\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{2},v\,(\boldsymbol{x})=-\frac{1}{16\pi},s\,(\boldsymbol{x})=\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{2}\ln\,(\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2}^{2})$. Assuming we did not know this exact solution, but at least knew $\Delta\,(u+sv)$ needed a logarithm singularity with the scaling and any non-singular terms undetermined, we could use $s\left(\boldsymbol{x}\right)=\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{2}\ln\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{2}$ and
@@
where $\{\boldsymbol{\omega}_{n}\}$ and $\{\tilde{\boldsymbol{\omega}}_{n}\}$ correspond to Fourier frequencies on the box $\Omega$. Note that $d_{n}$ and $\tilde{d}_{n}$ control the smoothness of $u$ and $v$, respectively.
It is important to note that $\Delta s,\partial_{x}s,$ and $\partial_{y}s$ should not be zero on $\Omega\setminus\left\{\boldsymbol{x}_{0}\right\}$ with a delta distribution singularity at $\boldsymbol{x}_{0}$; this would allow arbitrary multiples of $s,xs$ or $ys$ to be added to our solution and change the resulting integral (due to the delta distribution).
For our chosen $s$, a logarithm term appears in $\Delta s$, which will allow $\ln\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}$ to be integrated. After computing $u+sv$ so that $\Delta\left(u+sv\right)=-\frac{1}{2\pi}\ln\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}$ on $\tilde{N}$ scattered points, the boundary integral $\oint_{\partial S}\nabla\left(u+sv\right)\cdot\hat{\boldsymbol{n}}_{\partial S}$ is computed on 1000 evenly-spaced points on the unit circle with the standard trapezoidal rule (which converges super-algebraically for $C^{\infty}$ functions for the unit circle). The number of boundary points is chosen to ensure that the only significant component of the error is due to the surface discretization. Also note that if $\boldsymbol{x}_{0}$ is on or near $\partial S$, then a suitable quadrature for handling singular line integrals would be needed; a similar technique as the surface discretization or a more standard adaptive quadrature could be used.
We use $d_{n}=(\exp(q\sqrt{2\pi/T})+\exp(q\sqrt{\|\boldsymbol{\omega}_{n}\|_{2}}))^{2},q=4,T=10$, $\Omega=[-2,2]^{2}$, and $(2\cdot 30+1)^{2}$ Fourier basis functions. Interior points are generated using algorithm A.2 of [35] with 20 test points per point and no additional weighting of points near the boundary. The relative error for computing $\frac{1}{2\pi}\int_{S}\ln\|\boldsymbol{x}\|_{2}\,\text{d}\boldsymbol{x}=\frac{1}{4}$ is shown in Figure 4. High-order convergence is observed, where we note that the point spacing varies as $h_{\text{max}}\propto\tilde{N}^{-\frac{1}{2}}$. We note again that a unit circle is only used as a test problem to compare to a known solution; the method’s main use is for irregularly-shaped domains and surfaces.
@@ 4.2.2 Singular Integrand on a Surface
We now consider the surface described by $S:\{\left(x,y,z\right):\varphi\left(x,y,z\right)=0,z\geq 0\}$ where $\varphi$ is the level set for a paraboloid: $\varphi\left(x,y,z\right):=z+x^{2}+y^{2}-1.$ We compute
@@
where $\boldsymbol{x}_{0}=(0.6,0.4,0.48)\in S$. This integral can be interpreted as the electric potential at $\boldsymbol{x}_{0}$ produced by a constant surface charge on $S$. We integrate by attempting to find a function of the form $u+sv$ such that $\Delta_{S}\left(u+sv\right)=\frac{1}{4\pi}\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{-1}$, where $\Delta_{S}$ is now the Laplace-Beltrami operator, and then converting the surface integral to a line integral on the unit circle in the $xy$-plane. The line integral is again evaluated using the trapezoidal rule and 1000 evenly spaced points. The Laplace-Beltrami problem is discretized using a method similar to the one used in the previous test (see Eq. (37)), but with $\Delta_{S}$ computed using (31) replacing $\Delta$. We seek a solution of the form $u+sv$ where
(40) $s\left(\boldsymbol{x}\right)=\left\|\boldsymbol{x}-\boldsymbol{x}_{0}\right\|_{2},$
and $u,v\in C^{\infty}\left(\Omega\right)$ are unknown. This choice of $s$ gives $\Delta s=2\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{-1}$; however, $\Delta s$ is only one term in $\Delta_{S}s$. The necessary functions $u$ and $v$ to solve $\Delta_{S}\left(u+sv\right)=\frac{1}{4\pi}\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{-1}$ will vary depending on the local geometry of the surface and will be computed numerically. Whether such a solution will exist in the range of $\mathcal{E}$ is unknown, but the idea is simply to choose an $s$ term such that $\Delta_{S}s$ captures the singular behaviour in the Poisson problem. One can use (31) to show that $\Delta_{S}s-\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{-1}\to 0$ as $\boldsymbol{x}\to\boldsymbol{x}_{0}$ on $S$, noting that $((\boldsymbol{x}-\boldsymbol{x}_{0})/\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2})\cdot\hat{\boldsymbol{n}}_{S}=\mathcal{O}\,(\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2})$ as $\boldsymbol{x}\to\boldsymbol{x}_{0}$ for $\boldsymbol{x}$ on $S$. We will see that including the $s$ term drastically improves the numerical results.
The Hilbert space and $\mathcal{E}$ used are identical to Subsection 4.2.1, but with $\Omega=[-2,2]^{3}$ (instead of $2,2]^{2}$), $(2\cdot 11+1)^{3}$ Fourier basis functions, and $s$ given by (40). The singularity in the integrand is placed at $\boldsymbol{x}_{0}=(0.6,0.4,0.48)\in S$. Surface points are generated using version B of algorithm A.1 of [35] with 25 test points per point; notably, we do not place extra points near the singularity. This is compared to the naive approach, where a solution $u\in C^{\infty}(\Omega)$ is sought, in Table 5 (effectively searching for solutions of the form $\mathcal{E}\left(a,0\right)$ only, where $\mathcal{E}$ is similar to (38), but for 3 dimensions). We compare to a value computed (0.613817759) to machine precision from a polar integral resulting from a parametrization, and see that the augmented method converges to the first 7 or 8 digits by $\tilde{N}=2560$, while the naive approach fails to converge beyond a digit or two. The Poisson solution using the augmented approach for $\tilde{N}=2560$ is shown in Figure 5. We note that we do not increase the point density near the singularity at $\boldsymbol{x}_{0}$. We also point out that $\boldsymbol{x}_{0}$ is not at the vertex of $S$ and no boundary conditions are imposed, so the norm-minimizing solution need not be symmetric about the singularity or appear related to the boundary in any way.
@@
Figure 5: A computed solution to $\Delta_{S}u=\frac{1}{4\pi}\|\boldsymbol{x}-\boldsymbol{x}_{0}\|_{2}^{-1}$ on the paraboloid using the augmented method with $\tilde{N}=2560$ points ($\boldsymbol{x}_{0}=\left(0.6,0.4,0.48\right)$). The location of the singularity in the integrand is shown with a red circle.
@@ 5 Conclusions
We presented, analyzed, and tested two high-order, meshfree techniques for surface integration. Method 1 (Subsection 3.1) uses a result relating solvability and boundedness (Proposition 2.1) to determine which scaling of a function makes a Poisson problem solvable, then deduces the ratio of the integral of two functions on a domain of interest. This approach is useful for finding the average values of functions on surfaces, as shown in the tests of 4.1.1, and can be used directly on closed surfaces without splitting them into subdomains. We also demonstrated numerically that reasonable estimates of the average value of a function can be obtained using our approach even on irregular, non-uniform point distributions. Method 2 (Subsection 3.2) uses the divergence theorem to reduce surface integrals to line integrals and takes advantage of the ability of symmetric meshfree methods to find one of many possible solutions to PDEs; this avoids imposing boundary conditions. Avoiding boundary conditions is helpful, since PDEs with boundary conditions may not permit a smooth solution if the boundary is only piecewise smooth. Unlike method 1, method 2 computes integrals directly, not as a ratio of integrals. However, method 2 requires working on a surface with boundary; for closed surfaces we introduce an artificial boundary by cutting the surface into at least two patches, as discussed in 3.5. We demonstrated method 2’s accuracy in 4.1.2 and 4.1.3, where we estimated the surface area of a genus-two surface considerably more accurately than a standard triangulation approach. Both methods also preserve the super-algebraic approximation properties of meshfree interpolation methods and do not require prior knowledge of the probability distribution from which a point cloud is sampled; this differs from Monte Carlo methods which are $\mathcal{O}(\tilde{N}^{-\frac{1}{2}})$ and typically assume a certain sampling distribution.